IF ONE SIDE OFA RIGHT TRAINGLE IS 12 AND ITS HYPOTHENUSE IS 13, WHAT IS THE LENTH OF THE OTHER SIDE?
HELP!!!

Answer by aranajay(3)

let c is our hypotenous
let b is one side of right triangle
let a= is other side
Formula= c^2= a^2+ b^2
were looking for a?
so......a^2= c^2 - b^2
a^2= 13^2 - 12^2
A^2= 169- 144
A^2= 25
A = 5(Square root of 25 is 5) is the final answer